One of the Mars missions found a bacterial-like life form and NASA scientists identified the organism contains DNA as its genetic material, but in a slightly different form. On average the organism’s cells have 75 nuclei per cell, and the nuclear DNA is found to be around 4 billion bp in length. Using this information, calculate how many picograms of DNA each Martian cell has
Answer #1

Ans.) Answer = 306.75 picograms (pg)

By calculating molecular masses of different dNTPs and excluding molecular of water that is loss while forming one phosphodiester bond. We can calculate that

1 picogram = 978 Mbp = 0.978 × 109 bp

Given: Number of basepairs in each nucleus = 4 billions = 4 × 109

Total number of nuclei in each cell =75

Calculation:

Total number of bases in each martial cell = 75 × 4 × 109    = 300 × 109 bp

Total amount of DNA in picograms (pg) in each martial cell =

= 300 × 109 / 0.978 × 109 = [306.75 pg]